OpenStudy (anonymous):

How do you answer these questions? (operations with rational numbers) Can you also show me how to do it? 1. -2/5 - (-1/10 + 1/-2) 2.-3/5 (-3/4 - -1/4)

OpenStudy (anonymous):

\[-2/5-(-1/10 +1/-2)=-2/5-(-1/10-1/2)=-2/5-(-1/10-5/10)=-2/5-6/10=-4/10-6/10=-10/10=-1\]

OpenStudy (anonymous):

what about the second question?